• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Xiaofeng, Pibero

    Log entry time 13:24:15 on August18,2001

    Entry number 71161

    keyword=target NMR:43.99% EPR: 43.7%

    12:23 Last run numbers are 2902/21902
    12:23 Turn off transverse lasers
    12:24 Move target to Pick-Up Coils
    12:26 Rotate holding field from 270 to 0 deg
    12:30 Turn on longitudinal lasers
    12:36 NMR V=2.02 Vpp, P=43.99%

    12:46 EPR

    Conditions: B-field: 0 deg, lasers #5,6,7 on, grad coils off, pick-up coil position, beam off, temperature stable, last runs: 2902/21902

    P = 43.7%

    1st order fit:

    freq1: -47.6 kHz +/- 2.03 kHz | Polarisation: 44.51 % +/- 1.9 %
    freq2: 45.98 kHz +/- 1.88 kHz | Polarisation: 43 % +/- 1.76 %

    Comments: 1st order fit only since field was still drifting. Vmod = 0.85V

    12:51 Turn off longitudinal lasers
    12:52 Move target to Polarized He Cell position
    12:54 Rotate holding field from 0 to 270 deg